OpenStudy (anonymous):

Help! :/ Nineteen students play on a baseball team. How many ways can the batting order of 10 players be arranged?

OpenStudy (anonymous):

You want to find how many permutations of 10 there are from 19 people. So you calculate: 19P10 Do you know how to use this button on your calculator?

OpenStudy (anonymous):

I do not.

OpenStudy (anonymous):

If not, it's exactly: \[\frac{19!}{9!}\]

OpenStudy (anonymous):

How did you get that?

OpenStudy (anonymous):

The formula for nPr is: \[\frac{n!}{(n-r)!}\] nPr basically means "how many k permutations of n are there?" Does this make more sense? Also do you understand what I mean by factorial (!)?

OpenStudy (anonymous):

Yes and no

OpenStudy (anonymous):

Oh, no for the factorial? Factorial just means multiply that number by every positive number before it. So 9! means 9x8x7x6x5x4x3x2x1. Is this a homework question? If so I'm surprised that they'd give you this before explaining about factorials :P

OpenStudy (anonymous):

Yes I do online, this is just a practice but i just don't understand it, but that did help alot. So i get the 9x8x7x6x5x4x3x2x1 but how did you get 9?

OpenStudy (anonymous):

In fact I think I may have chosen the wrong one. It's nCr not nPr. The formula for nCr is: \[\frac{n!}{r!(n-r)!}\] Which is a lot smaller and looks like a more sensible answer. So for this question it'd be: \[\frac{19!}{10!\times 9!}\] You can work this out on a calculator or on wolfwram, it comes out as 92,378. Sorry about that :)

OpenStudy (anonymous):

nCr means "how many ways are there of choosing r from n".